OpenStudy (firejay5):

Use Synthetic Division. Show work on how you got your answer. 36.(2b^3 + b^2 - 2b + 3) / (b+1)^-1

OpenStudy (firejay5):

@ArkGoLucky Are you thinking???

OpenStudy (anonymous):

oh sorry I was doing something else

OpenStudy (anonymous):

this doesn't make sense \[\frac{1}{(b+1)^{-1}}=b+1\]

OpenStudy (firejay5):

Do you know how to do synthetic division??? @satellite73

OpenStudy (anonymous):

just multiply the big polynomial with b+1

OpenStudy (anonymous):

yes, but i cannot write it here i have tried, but it doesn't work it is just a technique, take a look at this and it explains well with an example http://www.purplemath.com/modules/synthdiv.htm

OpenStudy (firejay5):

@satellite73 Giving me a website won't help me

OpenStudy (anonymous):

(b+1)^-1 =1/(b+1) polynomial/(1/(b+1) = polynomial * (b+1)

OpenStudy (firejay5):

so it's 1/b+1

OpenStudy (firejay5):

@ArkGoLucky What goes in the box??? 

OpenStudy (anonymous):

that's not a division, the way it's written.

OpenStudy (firejay5):

I know that

OpenStudy (anonymous):

\[\frac{ 1 }{ (b+1)^{-1}} = b+1\]

OpenStudy (firejay5):

okay so do I right -1 in the box

OpenStudy (firejay5):

write

OpenStudy (anonymous):

what's 2/(2^-1) ?

OpenStudy (firejay5):

what are you doing

OpenStudy (anonymous):

2*2

OpenStudy (firejay5):

where did you get the 2

OpenStudy (anonymous):

1/(2^-1) = 2

OpenStudy (firejay5):

I mean it just doesn't look like you're helping me though, @Algebraic!

OpenStudy (anonymous):

this is not a division problem as written \[\frac{2b^3 + b^2 - 2b + 3}{(b+1)^{-1}}=(2b^3 + b^2 - 2b + 3) (b+1)\] is what @Algebraic! was telling you

OpenStudy (firejay5):

do you know what synthetic division is

OpenStudy (anonymous):

yes, but this is not a division problem, it is a multiplication problem

OpenStudy (firejay5):

exactly it's just called that

OpenStudy (firejay5):

so can you help me??? @satellite73 It's stumped me

OpenStudy (anonymous):

is the problem \[\frac{2b^3 + b^2 - 2b + 3}{b+1} \]??

OpenStudy (firejay5):

yea (2b^3 + b^2 - 2b + 3) / (b+1)^-1

OpenStudy (anonymous):

list the coefficients of the numerator, put a \(-1\) on the side 2 1 -2 3 -1 _________________

OpenStudy (anonymous):

bring down the first 2 2 1 -2 3 -1 _________________ 2

OpenStudy (anonymous):

then \(2\times -1=-2\) put it here 2 1 -2 3 -1 -2 _________________ 2

OpenStudy (anonymous):

then \(1-2=-1\) it goes below 2 1 -2 3 -1 -2 _________________ 2 -1

OpenStudy (anonymous):

then \(-1\times -1=1\) 2 1 -2 3 -1 -2 1 _________________ 2 -1

OpenStudy (anonymous):

\(-2+1=-1\) 2 1 -2 3 -1 -2 1 _________________ 2 -1 -1

OpenStudy (anonymous):

\(-1\times -1=1\) 2 1 -2 3 -1 -2 1 1 _________________ 2 -1 -1

OpenStudy (anonymous):

\(3+1=4\) 2 1 -2 3 -1 -2 1 1 _________________ 2 -1 -1 4 you are done

OpenStudy (firejay5):

not quite 2b^2 - b - 1 + 4/b+1

OpenStudy (anonymous):

polynomial if degree 3 divided by a polynomial of degree 1 is a polynomial of degree 2 your answer is \[2d^2-d-1+\frac{4}{d+1}\]

OpenStudy (firejay5):

it's b lol

OpenStudy (anonymous):

but at the risk of repeating what everyone is saying, as you wrote the problem it is NOT A DIVISION because the \((b+1)^{-1}\) in the denominator brings it up in to the numerator

OpenStudy (anonymous):

in other words \[\frac{2b^3 + b^2 - 2b + 3}{(b+1)^{-1}}=(2b^3 + b^2 - 2b + 3) (b+1)\]

OpenStudy (anonymous):

no division is indicated, it is a multiplication

OpenStudy (firejay5):

You know what the answer is for the long division is the same answer for this problem (synthetic division). Whatever you have leftover is put over the answer; for example: b +1, x-2, etc, so the answer is 4/b+1

OpenStudy (anonymous):

i guess there is no point in my repeating what i wrote above, so i will keep quiet

OpenStudy (firejay5):

so do you understand the answer is: 2b^2 - b - 1 + 4/ b+1

OpenStudy (firejay5):

never reciprocal it ----> b+1/4, because that's wrong

OpenStudy (anonymous):

You're missing the point @Firejay5 . Yes in the beginning it's division problem but if you simplify the expression, it becomes a multiplication problem.

OpenStudy (firejay5):

@ArkGoLucky It's over with and I am right; it's just another way of doing Long Division